Circular No. 8716

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2006cl, 2006cm, 2006cn, 2006co, 2006cp, 2006cq, 2006cr

Seven new apparent supernovae found on unfiltered CCD frames have been reported: 2006cl and 2006cn by the ‘Nearby Supernova Factory’ collaboration (cf. CBETs 520, 522); 2006cm and 2006cq by T. Puckett (with A. Langoussis for 2006cm, and with J. Newton and E. Briggs for 2006cq); 2006co by C. Colesanti, C. Jacques, E. Pimentel, and T. Napoleao (cf. IAUC 8658); 2006cp by H. Khandrika and W. Li (LOSS/KAIT; cf. IAUC 8710); and 2006cr by T. Boles (cf. IAUC 8699).

SN  2006 UT  α_{2000}  δ_{2000}  Mag.  Offset
2006cl May 20.4 16^{h}19^{m}48.^{s}80 −10^{o}03′25′′.2 19.2  –
2006cm May 24.37 21 20 17.46 − 1 41 02.7 16.6 2′′.0 W, 1′′.0 N
2006cn May 25.2 13 21 49.14 −23 51 04.6 19.0  –
2006co May 27.10 12 52 40.93 −39 01 48.9 16.0 24′′.4 E, 7′′.7 S
2006cp May 28.23 12 19 14.89 +22 25 38.2 18.2 19′′.9 E, 15′′.3 S
2006cq May 29.18 13 24 25.10 +30 57 22.7 17.3 4′′.9 W, 10′′.3 S
2006cr May 29.95 13 07 12.51 +58 08 06.1 18.1 8′′.6 W

Spectra show that 2006cl, 2006cn, and 2006co are type-II supernovae (details on CBETs 520, 522, and 529). Additional magnitudes for 2006cm in UGC 11723: 2005 Nov. 25, [19.3; 2006 May 25.37 UT, 16.4. A spectrum of 2006cm taken on May 29.47 showed it to be a reddened type-Ia supernova near maximum (details on CBET 526). Additional magnitudes for 2006co in ESO 323-G25: May 4.06, [18.5; 27.89, 15.9. Nothing was visible at the location of 2006co on a red Digitized Sky Survey (DSS) image from 1994. Additional magnitudes for 2006cp in UGC 7357: May 16.24, [19.0; 29.20, 17.8. A spectrum of 2006cp taken on May 30.26 showed it to be a type-Ia supernova around 1 week before maximum (details on CBET 528). Additional magnitudes for 2006cq in IC 4239: Jan. 20, [19.5 (Puckett); May 30.22, 17.3 (A. Sehgal, 0.50-m reflector, Osoyoos, BC). A spectrum of 2006cq obtained on May 31 showed it to be a type-Ia supernova close to maximum brightness (details on CBET 529). Additional magnitudes for 2006cr in UGC 8205: 1995 Apr. 23, [21.0 (DSS blue); 1997 June 25, [20.5 (DSS red); 2006 Apr. 5 and May 15, [19.5 (Boles). A spectrum of 2006cr taken on June 1 shows it to be a type-II supernova ∼ 1 week past explosion (details on CBET 531).

2006 June 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green